Blake Inc. needs €2,000,000 in 30 days. It can earn 5 percent annualized on a German security. The current spot rate for the euro is $1.00. Blake can borrow funds in the U.S. at an annualized interest rate of 6 percent. If Blake uses a money market hedge to hedge the payable, what is the cost of implementing the hedge?
Answer #1

Sol:

Amount needed by Blake Inc in 30 days = €2,000,000

Spot rate Dollar/Euro = $1

German rate for 30 days (Foreign currency) = 5% = 5% x 30/360 = 0.05 x 30/360 = 0.004167

U.S rate for 30 days (Home currency) = 6% = 6% x 30/360 = 0.06 x 30/360 = 0.005

Now invest amount receivable in Foreign currency = 2,000,000 / (1 + 0.004167) = 2,000,000 / 1.004167 = 1,991,700.5837

Now buy Foreign currency spot = 1,991,700.5837 x 1 = 1,991,700.5837

Now we need to borrow home currency ($) to buy foreign currency = 1,991,700.5837 x 1.005 = $2,001,659.0866

Therefore cost of implementing the hedge will be approximate $2,001,659.09